Solve the problem. Use the Central Limit Theorem.

In one region, the September energy consumption levels for single-family homes are found to be normally distributed with a mean of 1050 k Wh, and a standard deviation of 218 k Wh. If 50 different homes are randomly selected, find the probability that their mean energy consumption level for September is greater than 1075 k Wh.

Find the z-score (round to 2 places after decimal):

Is this an area to the left or to the right?

What is the area/probability?
Answer #1

P(\bar{X}>1075) = P\left ( \frac{\bar{X}-1050}{218/\sqrt{50}}>\frac{1075-1050}{218/\sqrt{50}} \right )

= P\left ( Z>0.81 \right )

= 1-P\left ( Z<0.81 \right )

We use excel function "NORMSDIST()" as :

= 1-P\left ( Z<0.811 \right )=1-NORMSDIST(0.811) = 1-0.7913 = 0.2087

1596042623729_image.png

This area is to the right